# Scattering wave packets of hadrons in gauge theories: Preparation on a quantum computer

### v2

@@ II.2 Ansatz for interacting wave-packet creation operator
The procedure to generate |$\ket{\Omega}$ will be described in Sec. III.1.1 for the example of the $Z_{2}$ LGT. The remainder of this section will be dedicated to the construction of $b_{k}^{\dagger}$ in each momentum sector, which eventually are assembled into the full wave-packet operator $b_{\Psi}^{\dagger}$.
The core of the interacting $b_{k}^{\dagger}$ is an ansatz proposed in Ref. Rigobello et al. (2021), in which wave packets of mesons in the lattice Schwinger model were built using tensor networks. This ansatz worked well only in the weak-coupling ($\epsilon\ll 1$) regime of the Schwinger model for system sizes studied in that work. As we will see, by adding a layer of classical or quantum optimization, this ansatz can be made suitable for both perturbative and non-perturbative regimes. To proceed, recall that in a confined gauge theory, $b_{k}^{\dagger}$ should only create gauge-invariant mesonic bound states. One can model $b_{k}^{\dagger}$ with a combination of mesonic creation operators consisting of fermion-antifermion creation-operator pairs:
@@
The form in Eq. (8) is motivated by the fact that for staggered free fermions (i.e., $U_{n}=1$ at all links), the expression in the square bracket in Eq. (8) reduces to $c_{p}^{\dagger}d_{q}^{\dagger}$, where $c_{p}^{\dagger}=\sum_{m}\mathscr{C}(p,m)\xi_{m}^{\dagger}$ and $d_{q}^{\dagger}=\sum_{n}\mathscr{D}(q,n)\xi_{n}$ diagonalize the free-fermion Hamiltonian in momentum space Rigobello et al. (2021). Note that by explicitly constructing the global $U(1)$ charge operator, one finds that the $c_{k}$-type excitations are the antiparticles of the $d_{k}$-type excitations (have an opposite U(1) charge). Similarly, by constructing the four-momentum operator, $P^{\mu}=(H,\bm{P})$, and demanding zero energy and momentum for such a state, it can be seen that $c_{k}$ and $d_{k}$ annihilate the vacuum Rigobello et al. (2021).
@@
The fermion-antifermion pairs, or bare mesons, are distributed in momentum space following the ansatz function $\eta(p,q)$. Similar to Ref. Rigobello et al. (2021), we consider a Gaussian distribution in relative momentum $p-q$:
@@
In Ref. Rigobello et al. (2021), $b_{k}^{\dagger}\ket{\Omega}$ describes the momentum eigenstate |$\ket{k}$ with $\sigma_{k}^{A}$ and $\mu_{k}^{A}$ manually tuned for each $k$. In this work, optimization on $\left(\sigma_{k}^{A},\mu_{k}^{A}\right)$ is explicitly performed by searching for the lowest-energy state with $b_{k}^{\dagger}$ excitations in each $k$ sector. For small systems, $b_{k}^{\dagger}$ with the optimized parameters is benchmarked against exact-diagonalization results to ensure that |$\ket{k}$ is indeed created as desired. The optimization strategy and results will be discussed thoroughly in the next section. For larger systems, one can resort to a variational quantum eigensolver (VQE) to perform energy minimization in each sector using a quantum computer. Such details will be presented in Sec. III.
@@
Here, ~$\widetilde{\mathcal{M}}_{m,n}$ is the Jordan-Wigner transformed $\mathcal{M}_{m,n}$ that is obtained by substituting Eq. (II.1) in the expression for $\mathcal{M}_{m,n}$ given below Eq. (8). For example, consider $m<n$ and $n-m<N/2$, then a forward-wrapped meson creation operator $\mathcal{M}_{m,n}$ leads to
@@
$C_{m,n}=\sum_{k}\Psi(k)\sum_{p,q\in\widetilde{\Gamma}}\delta_{k,p+q}\eta(p,q)\sqrt{\frac{m_{f}+\omega_{p}}{2\pi\omega_{p}}}\sqrt{\frac{m_{f}+\omega_{q}}{2\pi\omega_{q}}}{e^{i(pm+qn)}}\left(\mathcal{P}_{m0}+v_{p}\mathcal{P}_{m1}\right)\left(-v_{q}\mathcal{P}_{n0}+\mathcal{P}_{n1}\right).$ (14)
@@
We end this section with a remark. The wave-packet creation operator in Eq. (12) can be generalized to non-Abelian gauge theories in 1+1 D. Let us consider $SU(N)$ LGTs in which fermions transform in the fundamental representation of the gauge group, as in the Standard Model, and further consider only one flavor of fermions as in this work. Then, the only changes in the ansatz are to take into account the multi-component nature of the fermions (hence a generalized Jordan-Wigner transform needs to be applied Davoudi et al. (2022b); Raychowdhury and Stryker (2020)) and to account for the multi-component nature of link operators that are now non-commuting matrices, hence their order of operation in the ansatz matters. Mesons are still states that carry no gauge-group charges (i.e., colors). There will be more possibilities for building mesons using different components (colors) of fermions matched to the appropriate components (color indices) of the gauge links. Nonetheless, the general form of the ansatz in this section should still apply, although a more elaborate form for the internal wave-function profiles may be needed. We leave developing a detailed ansatz for mesonic excitations in non-Abelian theories in 1+1 D to future work, and here we proceed with testing the fidelity of such an ansatz for the case of the $Z_{2}$ and $U(1)$ LGTs in 1+1 D. Outlook for generalizing the ansatz to higher-dimensional theories and other gauge groups is further discussed in Sec. IV.
II.3 Numerical verification of the ansatz in small systems
@@
In this section, we benchmark the wave-packet ansatz against exact results for both the $Z_{2}$ and $U(1)$ LGTs coupled to one flavor of staggered fermions in 1+1 D. Our hardware results for wave-packet preparation in the next section are limited to the case of a 6-site theory due to computational cost. Therefore, we focus on demonstrating the goodness of the optimized ansatz for the same system size in this section. Nonetheless, we have verified that the ansatz works well for larger systems and a range of model parameters, although the fidelity tends to drop for momenta near the edges of the Brillouin zone. Examples of this numerical benchmark for the case of a 10-site theory are provided in Appendix A.
@@
Figure 5: The $k$-optimization heat maps for the 6-site theories. Panels (a) and (b) refer to the $Z_{2}$ LGT and the $U(1)$ LGT (with a cutoff of $\Lambda=1$), respectively. The region left to the cyan contour is where the non-mesonic excitations are less energetic than the single-meson |$\ket{k}=0$ state. The results for |$\ket{k=-\frac{\pi}{3}}$ are identical to those for |$\ket{k=\frac{\pi}{3}}$ and are not shown.
@@
For a lattice with 6 staggered sites, the Brillouin zone contains 3 discrete momentum sectors, $k\in\widetilde{\Gamma}=\{-\frac{\pi}{3},0,\frac{\pi}{3}\}$. So one needs to optimize the ansatz for $b_{k}^{\dagger}$ for these $k$ values. First, we examine whether high fidelities can be achieved for $k$-momentum-state optimization with different model parameters. The fidelity of each $k$-momentum state is defined as:
@@
where the subscript exact stands for exact-diagonalized momentum eigenstate |$\ket{k}$. The fidelity results of scanning $(m_{f},\epsilon)$ parameter pairs are shown in Fig. 5. The optimization performs best in the strong-coupling regime for both theories, and is able to reach a fidelity $>0.95$ in the intermediate coupling regime where the energy partition for the hopping term is comparable to the mass term and the electric-field term. The parameters $m_{f}=1$ and $\epsilon=-0.3$ in the $Z_{2}$ LGT, used in the hardware demonstration in the next section, correspond to $F>0.98$ for all $k$.
@@
Figure 6: Optimization results for $k=0~{}(\frac{\pi}{3})$ momentum eigenstates are shown in the left (right) panel for the case of $Z_{2}$ LGT with $m_{f}=1$ and $\epsilon=0.3$ (top) and $U(1)$ LGT with $m_{f}=1$ and $\epsilon=1$ (bottom). The physical basis-state probabilities the of $k=-\frac{\pi}{3}$ state are identical to $k=\frac{\pi}{3}$ and are not shown. The physical basis states are listed in Tables 2 and 3 in Appendix E.
@@
The ansatz creation operator is capable of creating mesonic excitations on top of the interacting ground state |$\ket{\Omega}$. However, not all states in the interacting spectrum can be generated by such operators. Under PBCs, there are other non-mesonic excitations generated by operators made up of only the gauge-link operators spanning over the entire lattice, e.g., $\prod_{n=0}^{N-1}U_{n}$ and its conjugate. As a result, the energies of these non-mesonic excitations grow with $|\epsilon|$ and the system size $N$. Furthermore, excitations consisting of only these non-mesonic operators have momentum $k=0$ since these operators are manifestly invariant under translation. In large lattices, states generated from non-mesonic excitations have larger energy compared to the mesonic excitations captured by the ansatz of this work. However, for smaller systems and sufficiently small $|\epsilon|$ values, the $k=0$ sector contains states generated by purely non-mesonic operators that have energies comparable or smaller than the states consisting of purely mesonic excitations. Such a transition in the Hamiltonian parameter space for 6-site theories is denoted by the cyan contour for the $k=0$ heat maps in Fig. 5. In the region left (right) of the cyan contour, the first excited state is created by the purely non-mesonic (mesonic) operators. Importantly, our ansatz captures the mesonic states with good fidelity in both regions.
@@
Second, choosing the parameters $m_{f}=1$ and $\epsilon=0.3$ as in the next section for the case of $Z_{2}$ and $m_{f}=1$ and $\epsilon=1$ for the case of $U(1)$, we plot in Fig. 6 the probability distribution of each of the physical basis states (called ‘physical configurations’ throughout) in both theories, comparing those obtained from the optimized ansatz and exact diagonalization. It is seen that each $k$-momentum state created by the ansatz recovers the exact-diagonalized state probabilities.
III Preparing mesonic wave packets in $Z_{2}$ LGT in 1+1 D
@@
III.1 Algorithm and circuit design
Our wave-packet preparation algorithm, as summarized in the right panel of Fig. 2, involves four steps: the ground(vacuum)-state preparation, the $k$-momentum-eigenstate optimization, and the wave-packet assembly and circuit implementation. This section details the algorithm for the above components and demonstrates the design of the associated quantum circuits. An important aspect of the circuit is to limit the resources required to fit into Noisy Intermediate-Scale Quantum (NISQ) hardware. With this in mind, we prepare the $Z_{2}$ interacting ground state, |$\ket{\Omega}$, and excited states, |$\ket{k}$, based on a VQE, a class of hybrid classical-quantum algorithms suitable especially for the NISQ era of computing Peruzzo et al. (2014); McClean et al. (2016); Tilly et al. (2022). Applications of VQE-based algorithms to gauge theories have been extensively explored in recent years for ground-state properties and dynamics Kokail et al. (2019); Lumia et al. (2022); Farrell et al. (2023b); Paulson et al. (2021); Atas et al. (2021); Popov et al. (2023); Liu et al. (2022); Nagano et al. (2023); Xie et al. (2022). Once |$\ket{\Omega}$ and |$\ket{k}$ are obtained, a wave-packet preparation circuit can be derived and attached to the ground-state preparation circuit. We list the complete state-preparation algorithm below:
@@ III.1 Algorithm and circuit design
Figure 7: The schematic diagram of the full algorithm. The numbers correspond to the steps described in Sec. III.1. The purple block in step 1 depicts the circuit for $\mathcal{Q}_{\rm GS}$ parameterized by $\left(\theta^{h},\theta^{\epsilon}\right)$ with $N_{\rm GS}=1$, see Eq. (17). The evolution operators corresponding to the hopping (electric-field) terms in the Hamiltonian are denoted by multi-(single-)qubit gates. Qubits representing fermions and bosons are labeled by $f$ and $b$, respectively, and the subscripts are the site indices. The qubit mapping of the initial state |$\ket{\Omega}_{0}$ is shown next to its basis-state representation. VQE is applied to find the optimum parameters $\left(\theta^{h*},\theta^{\epsilon*}\right)$ that minimize the energy with respect to the $Z_{2}$ LGT Hamiltonian. The blue block in the top (bottom) circuit in step 2 denotes the circuit for $\mathcal{Q}_{k}$ parameterized by the ansatz parameters $(\sigma_{k}^{A},\mu_{k}^{A})$, for $k=\pm\frac{\pi}{3}$ ($k=0$). VQE is employed to obtain the parameters $(\sigma_{k}^{A*},\mu_{k}^{A*})$ that optimize the ansatz for each $k$. The full circuit that prepares a wave packet |$\ket{\Psi}$ from |$\ket{\Omega}_{0}$ is shown in step 4, with the ancilla qubit labeled by $a$. The repeated blue blocks acting on |$\ket{\Omega}$ denote the $\mathcal{Q}_{\rm WP}(C_{m,n})$ circuit with each blue block implementing one Trotter step in the Trotterized Eq. (19). The green block indicates the SVD circuit for $e^{-i\theta\Theta_{m,n}}$ and is described in detail in Fig. 8.
@@
III.1.1 Ground-state preparation
Here, we demonstrate a VQE-based ground-state preparation method that meets the requirement of a shallow circuit with low gate counts. We follow the general strategy in Ref. Lumia et al. (2022) where the desired entanglement of the ground state is built in by evolving some initial state with operators that respect the symmetry of the theory. The natural candidates for the evolution operators are, therefore, the hopping term, the mass term, and the electric-field term in the Hamiltonian (labeled $H^{h}=\sum_{n\in\Gamma}H_{n,n+1}$, $H^{m}=\sum_{n\in\Gamma}H^{m}_{n}$, and $H^{\epsilon}=\sum_{n\in\Gamma}H^{\epsilon}_{n}$ below, respectively). The initial state |$\ket{\psi}_{0}$ is chosen to be the strong-coupling vacuum |$\ket{\Omega}_{0}$. Explicitly,

### v3

@@ 2.2 Ansatz for interacting wave-packet creation operator
The procedure to generate |$\ket{\Omega}$ will be described in Sec. 3.1.1 for the example of the $Z_{2}$ LGT. The remainder of this section will be dedicated to the construction of $b_{k}^{\dagger}$ in each momentum sector, which eventually are assembled into the full wave-packet operator $b_{\Psi}^{\dagger}$.
The core of the interacting $b_{k}^{\dagger}$ is an ansatz proposed in Ref. [33], in which wave packets of mesons in the lattice Schwinger model were built using tensor networks. This ansatz worked well only in the weak-coupling ($\epsilon\ll 1$) regime of the Schwinger model for system sizes studied in that work. As we will see, by adding a layer of classical or quantum optimization, this ansatz can be made suitable for both perturbative and non-perturbative regimes. To proceed, recall that in a confined gauge theory, $b_{k}^{\dagger}$ should only create gauge-invariant mesonic bound states. One can model $b_{k}^{\dagger}$ with a combination of mesonic creation operators consisting of fermion-antifermion creation-operator pairs:
@@
The form in Eq. (8) is motivated by the fact that for staggered free fermions (i.e., $U_{n}=1$ at all links), the expression in the square bracket in Eq. (8) reduces to $c_{p}^{\dagger}d_{q}^{\dagger}$, where $c_{p}^{\dagger}=\sum_{m}\mathscr{C}(p,m)\xi_{m}^{\dagger}$ and $d_{q}^{\dagger}=\sum_{n}\mathscr{D}(q,n)\xi_{n}$ diagonalize the free-fermion Hamiltonian in momentum space [33]. Note that by explicitly constructing the global $U(1)$ charge operator, one finds that the $c_{k}$-type excitations are the antiparticles of the $d_{k}$-type excitations (have an opposite U(1) charge). Similarly, by constructing the four-momentum operator, $P^{\mu}=(H,\bm{P})$, and demanding zero energy and momentum for such a state, it can be seen that $c_{k}$ and $d_{k}$ annihilate the vacuum [33].
@@
The fermion-antifermion pairs, or bare mesons, are distributed in momentum space following the ansatz function $\eta(p,q)$. Similar to Ref. [33], we consider a Gaussian distribution in relative momentum $p-q$:
@@
In Ref. [33], $b_{k}^{\dagger}\ket{\Omega}$ describes the momentum eigenstate |$\ket{k}$ with $\sigma_{k}^{A}$ and $\mu_{k}^{A}$ manually tuned for each $k$. In this work, optimization on $\left(\sigma_{k}^{A},\mu_{k}^{A}\right)$ is explicitly performed by searching for the lowest-energy state with $b_{k}^{\dagger}$ excitations in each $k$ sector. For small systems, $b_{k}^{\dagger}$ with the optimized parameters is benchmarked against exact-diagonalization results to ensure that |$\ket{k}$ is indeed created as desired. The optimization strategy and results will be discussed thoroughly in the next section. For larger systems, one can resort to a variational quantum eigensolver (VQE) to perform energy minimization in each sector using a quantum computer. Such details will be presented in Sec. 3.
@@
Here, ~$\widetilde{\mathcal{M}}_{m,n}$ is the Jordan-Wigner transformed $\mathcal{M}_{m,n}$ that is obtained by substituting Eq. (3) in the expression for $\mathcal{M}_{m,n}$ given below Eq. (8). For example, consider $m<n$ and $n-m<N/2$, then a forward-wrapped meson creation operator $\mathcal{M}_{m,n}$ leads to
@@
$C_{m,n}=\sum_{k}\Psi(k)\sum_{p,q\in\widetilde{\Gamma}}\delta_{k,p+q}$ $\eta(p,q)\sqrt{\frac{m_{f}+\omega_{p}}{2\pi\omega_{p}}}\sqrt{\frac{m_{f}+\omega_{q}}{2\pi\omega_{q}}}$
×$\times{e^{i(pm+qn)}}\left(\mathcal{P}_{m0}+v_{p}\mathcal{P}_{m1}\right)\left(-v_{q}\mathcal{P}_{n0}+\mathcal{P}_{n1}\right).$ (14)
@@
We end this section with a remark. The wave-packet creation operator in Eq. (12) can be generalized to non-Abelian gauge theories in 1+1 D. Let us consider $SU(N)$ LGTs in which fermions transform in the fundamental representation of the gauge group, as in the Standard Model, and further consider only one flavor of fermions as in this work. Then, the only changes in the ansatz are to take into account the multi-component nature of the fermions (hence a generalized Jordan-Wigner transform needs to be applied [54, 122]) and to account for the multi-component nature of link operators that are now non-commuting matrices, hence their order of operation in the ansatz matters. Mesons are still states that carry no gauge-group charges (i.e., colors). There will be more possibilities for building mesons using different components (colors) of fermions matched to the appropriate components (color indices) of the gauge links. Nonetheless, the general form of the ansatz in this section should still apply, although a more elaborate form for the internal wave-function profiles may be needed. We leave developing a detailed ansatz for mesonic excitations in non-Abelian theories in 1+1 D to future work, and here we proceed with testing the fidelity of such an ansatz for the case of the $Z_{2}$ LGT in 1+1 D. To keep the discussion concise, similar fidelity tests for the $U(1)$ LGT are presented in Appendix A. Outlook for generalizing the ansatz to higher-dimensional theories and other gauge groups is further discussed in Sec. 4.
2.3 Numerical verification of the ansatz in small systems
@@
In this section, we benchmark the wave-packet ansatz against exact results for the $Z_{2}$ LGT coupled to one flavor of staggered fermions in 1+1 D. Our hardware results for wave-packet preparation in the next section are limited to the case of a 6-site theory due to computational cost. Therefore, we focus on demonstrating the goodness of the optimized ansatz for small systems (6 and 10 sites) in this section. For the $U(1)$ LGT, verification of the ansatz’s performance, similarly for a range of model parameters, is presented in Appendix A.
@@
Figure 5: The $k$-optimization heat maps for the $Z_{2}$ LGT. Panels (a) and (b) refer to the 6-site and the 10-site theories, respectively. The region left to the cyan contour is where the non-mesonic excitations are less energetic than the single-meson |$\ket{k}=0$ state. For the 6-site theory, the results for |$\ket{k=-\frac{\pi}{3}}$ are identical to those for |$\ket{k=\frac{\pi}{3}}$ and are not shown. Similarly, the optimization results for $k$ and $-k$ values are identical in the 10-site theory and only positive momentum eigenstates are presented. The ansatz works better in the strong coupling regime, and for momenta closer to the center of the Brillouin zone ($k=0$).
@@
For a lattice with 6 (10) staggered sites, the Brillouin zone contains 3 (5) discrete momentum sectors, $k\in\widetilde{\Gamma}=\{-\frac{\pi}{3},0,\frac{\pi}{3}\}(\widetilde{\Gamma}=\{-\frac{2\pi}{5},-\frac{\pi}{5},0,\frac{\pi}{5},\frac{2\pi}{5}\})$. So one needs to optimize the ansatz for $b_{k}^{\dagger}$ for these $k$ values. First, we examine whether high fidelities can be achieved for $k$-momentum-state optimization with different model parameters. The fidelity of each $k$-momentum state is defined as:
@@
where the subscript exact stands for exact-diagonalized momentum eigenstate |$\ket{k}$. The fidelity results of scanning $(m_{f},\epsilon)$ parameter pairs are shown in Fig. 5. The optimization performs best in the strong-coupling regime for both lattices. It is also able to reach a fidelity $>0.95$ in the intermediate coupling regime (where the energy partition for the hopping term is comparable to the mass term and the electric-field term), except for $k=\pm\frac{2\pi}{5}$ in the 10-site theory. There, the fidelity between the optimized state and the exact-diagonalized state is around 0.90 for most of the parameter space. It is generally observed that the $k$-momentum optimization becomes less accurate with $k$ closer to the boundary of the Brillouin zone, $\widetilde{\Gamma}$. Therefore, if the wave packet is centered further away from the edge of the Brillouin zone, the Gaussian profile of the wave packet will suppress the effect of these lower-fidelity high-momentum states in the full wave-packet state. The parameters $m_{f}=1$ and $\epsilon=-0.3$ in the 6-site $Z_{2}$ LGT, used in the hardware demonstration in the next section, correspond to $F>0.98$ for all $k$.
@@
Figure 6: Optimization results for $k=0\leavevmode\nobreak\ (\frac{\pi}{3})$ momentum eigenstates are shown in the left (right) panel for the case of the 6-site $Z_{2}$ LGT with $m_{f}=1$ and $\epsilon=-0.3$. The physical basis-state probabilities the of $k=-\frac{\pi}{3}$ state are identical to $k=\frac{\pi}{3}$ and are not shown. The physical basis states of the 6-site $Z_{2}$ theory are listed in Tables 2.
@@
The ansatz creation operator is capable of creating mesonic excitations on top of the interacting ground state |$\ket{\Omega}$. However, not all states in the interacting spectrum can be generated by such operators. Under PBCs, there are other non-mesonic excitations generated by operators made up of only the gauge-link operators spanning over the entire lattice, e.g., $\prod_{n=0}^{N-1}U_{n}$ and its conjugate. As a result, the energies of these non-mesonic excitations grow with $|\epsilon|$ and the system size $N$. Furthermore, excitations consisting of only these non-mesonic operators have momentum $k=0$ since these operators are manifestly invariant under translation. In large lattices, states generated from non-mesonic excitations have larger energy compared to the mesonic excitations captured by the ansatz of this work. However, for smaller systems and sufficiently small $|\epsilon|$ values, the $k=0$ sector contains states generated by purely non-mesonic operators that have energies comparable or smaller than the states consisting of purely mesonic excitations. Such a transition in the Hamiltonian parameter space is denoted by the cyan contour for the $k=0$ heat maps in Fig. 5. In the region left (right) of the cyan contour, the first excited state is created by the purely non-mesonic (mesonic) operators. As expected, this region in the 10-site theory is smaller than that in the 6-site theory. Importantly, our ansatz captures the mesonic states with good fidelity in both regions.
@@
Second, choosing the parameters $m_{f}=1$ and $\epsilon=-0.3$ as in the next section for the case of $Z_{2}$, we plot in Fig. 6 the probability distribution of each of the physical basis states (called ‘physical configurations’ throughout) in the $Z_{2}$ theory, comparing those obtained from the optimized ansatz and exact diagonalization. It is seen that each $k$-momentum state created by the ansatz recovers the exact-diagonalized state probabilities.
3 Preparing mesonic wave packets in $Z_{2}$ LGT in 1+1 D
@@
3.1 Algorithm and circuit design
Our wave-packet preparation algorithm, as summarized in the right panel of Fig. 2, involves four steps: the ground(vacuum)-state preparation, the $k$-momentum-eigenstate optimization, and the wave-packet assembly and circuit implementation. This section details the algorithm for the above components and demonstrates the design of the associated quantum circuits. An important aspect of the circuit is to limit the resources required to fit into Noisy Intermediate-Scale Quantum (NISQ) hardware. With this in mind, we prepare the $Z_{2}$ interacting ground state, |$\ket{\Omega}$, and excited states, |$\ket{k}$, based on a VQE, a class of hybrid classical-quantum algorithms suitable especially for the NISQ era of computing [105, 106, 107]. Applications of VQE-based algorithms to gauge theories have been extensively explored in recent years for ground-state properties and dynamics [123, 124, 125, 57, 52, 95, 109, 126, 127]. Once |$\ket{\Omega}$ and |$\ket{k}$ are obtained, a wave-packet preparation circuit can be derived and attached to the ground-state preparation circuit. We list the complete state-preparation algorithm below:
@@ 3.1 Algorithm and circuit design
Figure 7: The schematic diagram of the full algorithm. The numbers correspond to the steps described in Sec. 3.1. The purple block in step 1 depicts the circuit for $\mathcal{Q}_{\rm GS}$ parameterized by $\left(\theta^{h},\theta^{\epsilon}\right)$ with $N_{\rm GS}=1$, see Eq. (17). The evolution operators corresponding to the hopping (electric-field) terms in the Hamiltonian are denoted by multi-(single-)qubit gates. Qubits representing fermions and bosons are labeled by $f$ and $b$, respectively, and the subscripts are the site indices. The qubit mapping of the initial state |$\ket{\Omega}_{0}$ is shown next to its basis-state representation. VQE is applied to find the optimum parameters $\left(\theta^{h*},\theta^{\epsilon*}\right)$ that minimize the energy with respect to the $Z_{2}$ LGT Hamiltonian. The blue block in the top (bottom) circuit in step 2 denotes the circuit for $\mathcal{Q}_{k}$ parameterized by the ansatz parameters $(\sigma_{k}^{A},\mu_{k}^{A})$, for $k=\pm\frac{\pi}{3}$ ($k=0$). VQE is employed to obtain the parameters $(\sigma_{k}^{A*},\mu_{k}^{A*})$ that optimize the ansatz for each $k$. The full circuit that prepares a wave packet |$\ket{\Psi}$ from |$\ket{\Omega}_{0}$ is shown in step 4, with the ancilla qubit labeled by $a$. The repeated blue blocks acting on |$\ket{\Omega}$ denote the $\mathcal{Q}_{\rm WP}(C_{m,n})$ circuit with each blue block implementing one Trotter step in the Trotterized Eq. (19). The green block indicates the SVD circuit for $e^{-i\theta\Theta_{m,n}}$ and is described in detail in Fig. 8.
@@
3.1.1 Ground-state preparation
Here, we demonstrate a VQE-based ground-state preparation method that meets the requirement of a shallow circuit with low gate counts. We follow the general strategy in Ref. [124] where the desired entanglement of the ground state is built in by evolving some initial state with operators that respect the symmetry of the theory. The natural candidates for the evolution operators are, therefore, the hopping term, the mass term, and the electric-field term in the Hamiltonian (labeled $H^{h}=\sum_{n\in\Gamma}H_{n,n+1}$, $H^{m}=\sum_{n\in\Gamma}H^{m}_{n}$, and $H^{\epsilon}=\sum_{n\in\Gamma}H^{\epsilon}_{n}$ below, respectively). The initial state |$\ket{\psi}_{0}$ is chosen to be the strong-coupling vacuum |$\ket{\Omega}_{0}$. Explicitly,